# ARFlow: Human Action-Reaction Flow Matching with Physical Guidance

### v2

@@ 4 Experiments
Our experiment setting of human action-reaction synthesis is online, unconstrained as in  [69] for its significant potential for practical applications. Online represents real-time reaction generation where future motions of the actor are not visible to the reactor, and the opposite is offline to relax the synchronicity. Unconstrained means that the intention of the actor is invisible to the reactor. To demonstrate the universality of our flow matching and physical guidance methods, we also conducted offline setting experiments.
@@
ReGenNet $\underline{11.00^{\pm 0.74}}$ $\mathbf{0.749^{\pm 0.0002}}$ $13.80^{\pm 0.16}$ $\underline{22.90^{\pm 0.14}}$
@@
ReGenNet $\underline{13.76^{\pm 4.78}}$ $0.601^{\pm 0.0040}$ $6.35^{\pm 0.24}$ $\underline{12.02^{\pm 0.33}}$
@@
ReGenNet $\underline{6.19^{\pm 0.33}}$ $0.772^{\pm 0.0003}$ $14.03^{\pm 0.09}$ $\underline{25.21^{\pm 0.34}}$
@@
Table 3: Results on the offline, unconstrained setting on NTU120-AS. We highlight the best result in Bold and the second best in underline.
FID $\downarrow$ IF $\downarrow$ IV $\downarrow$
Real 0.09 38.38% 3.33
       Offline
MDM 22.04 29.77% 2.50
ReGenNet 6.19 35.77% 2.85
ARFlow 5.00 23.91% 2.01
MDM ${}^{\text{guidance}}$ 20.42 25.58% 2.10
ReGenNet ${}^{\text{guidance}}$ 7.69 25.10% 1.92
ARFlow ${}^{\text{guidance}}$ 5.19 15.82% 1.31
       Online
MDM 51.11 51.87% 4.22
ReGenNet 11.00 31.92% 2.63
ARFlow 7.89 17.44% 1.55
MDM ${}^{\text{guidance}}$ 40.03 44.84% 3.67
ReGenNet ${}^{\text{guidance}}$ 11.63 17.11% 1.37
ARFlow ${}^{\text{guidance}}$ 8.75 8.56% 0.76
[…]
Guidance w. $\mathcal{L}_{\text{pene}}$ $8.751^{\pm 0.2047}$ $0.740^{\pm 0.0002}$ $13.613^{\pm 0.1004}$ $24.096^{\pm 0.1433}$
@@
Guidance w. $\mathcal{L}_{\text{pene}}$ $5.196^{\pm 0.1144}$ $0.753^{\pm 0.0002}$ $13.824^{\pm 0.0839}$ $25.164^{\pm 0.1630}$
@@
Table 6: Ablation studies on the offline, unconstrained setting on the NTU120-AS dataset. Bold indicates the best result in our method.
4.1 Experiment setup
Datasets. We evaluate our model on NTU120-AS and Chi3D-AS datasets with SMPL-X [46] body models and actor-reactor annotations as in [47].They contains 8118 and 373 human interaction sequences, respectively. “AS”[69] represents that they are an extended version of the original dataset [17, 40, 61], which adds annotations to distinguish actor-reactor order of each interaction sequence and SMPL-X body models for more detailed representations. We adopt the 6D rotation representation [79] in all our experiments.
4.2 Comparison to baselines
For the NTU120-AS dataset in Tab. 1 and Chi3D-AS dataset in Tab. 2, our proposed ARFlow notably outperforms baselines in terms of the FID metric, demonstrating that our method better models the mapping between the action and the reaction distribution. Our method achieves the best FID and multi-modality, second best for the action recognition accuracy and diversity on the NTU120-AS dataset and the best FID and multi-modality, second best for the diversity on the Chi3D-AS dataset. For a fair comparison, we use the pre-trained action recognition model in [69], so our action recognition accuracy is very close to the results of [69]. Given the restricted size of the Chi3D-AS test set, some fluctuations in the experimental results are to be expected.
As for the offline setting, we replace the Transformer decoder units equipped with attention masks with an 8-layer Transformer encoder architecture just like ReGenNet [69]. As demonstrated in Tab. 3, our model achieves superior performance on most of the metrics, highlighting the effectiveness and versatility of our method across various scenarios.
4.3 Physical Constraint Guidance
Qualitative results. In Fig. 3, we provide a qualitative comparison of reaction sequences generated by MDM, ReGenNet and our ARFlow with physical constraint guidance, for the same action sequences sampled from the test sets of NTU120-AS datasets. The visualization results demonstrate that both MDM and ReGenNet produce varying degrees of penetration between the actor and the reactor. In contrast, our method not only produces more physically plausible reactions, i.e., mitigate penetration problems, but also more responsive reactions, i.e., move backward when kicked, due to stronger modeling ability for causal relationship between actions and reactions. For more visualizations and videos of the generated human reactions, please refer to the supplementary materials.
Quantitative results. In Tab. 4, our ARFlow with physical constraint guidance achieves the lowest Intersection Volume, Intersection Frequency and FID than other baselines, which shows that our method achieves the lowest level of penetration while ensuring the highest generation quality. In addition, other baseline methods using our physical constraint guidance significantly reduce the degree of penetration, which demonstrates the effectiveness and universality of our guidance method. Although our guidance method effectively suppresses penetration, it also leads to a slight increase in FID, as FID only measures the similarity between generated results and the ground truth distribution and the dataset itself exists a certain degree of penetration.
4.4 Ablation Study
Network Prediction. As depicted in Sec. 3.1, a straightforward and effective strategy is to estimate clean body poses directly through a neural network, i.e., $\mathbf{x}_{1}$-prediction. We compared it with $\mathbf{v}$-prediction and the results are listed on the Prediction setting in Tab. 5 and  Tab. 6. Obviously, $\mathbf{x}_{1}$-prediction has demonstrated superior performance across both online and offline settings. The reason we analyze it is that the geometric losses to regularize the generative network during the training phase directly acts on the predicted clean body poses, while $\mathbf{v}$-prediction requires using the predicted vector field to estimate the clean poses through Equation 13, so the models trained by $\mathbf{x}_{1}$-prediction will achieve stronger overall performance in the task of human action-reaction synthesis.
Physical constraint guidance. As we discussed earlier, the results in Tab. 4 indicate that our physical constraint guidance can effectively suppress penetration occurrences, and it may also lead to a decrease in some other metrics shown in Tab. 5 and Tab. 6. We also provide a qualitative comparison of the effects before and after using our physical constraint guidance in Fig. 10.1. The qualitative and quantitative results demonstrated that our method achieves the lowest penetration level while maintaining the best quality of generated reactions.
Number of Euler sampling timesteps. We present comprehensive evaluation results in both online and offline scenarios, with varying Euler sampling intervals (2, 5, 10 and 100 timesteps), including the latency of reaction generation per frame on online settings and overall latency on offline settings. The experimental results, as detailed in Tab. 5 and Tab. 6, suggest that the 5-timestep Euler sampling consistently achieves optimal performance, demonstrating superior FID scores while maintaining low latency across both evaluation settings. Thus, we adopt the 5-timestep inference as the standard configuration like  [69] for all the experimental results reported in this study.
5 Conclusion
In this work, we have presented Action-Reaction Flow Matching (ARFlow), a novel framework for human action-reaction synthesis that addresses the limitations of existing diffusion-based approaches. By establishing direct action-to-reaction mappings through flow matching, ARFlow eliminates the need for complex conditional mechanisms and computationally expensive noise-to-data transformations. ARFlow involves a $\mathbf{x}_{1}$-prediction sampling algorithm with physical constraint guidance to enable efficient, physically plausible motion generation while preventing body penetration artifacts. Extensive evaluations on the NTU120 and Chi3D datasets demonstrate that ARFlow excels over existing methods, showing superior performance in terms of Fréchet Inception Distance and motion diversity. Additionally, it significantly reduces body collisions, as evidenced by our new Intersection Volume and Intersection Frequency metrics.
However, while ARFlow’s physical constraint guidance proves effective in reducing body penetrations, it may lead to manifold distortions—deviations from the natural motion distribution established by flow matching. Addressing this challenge opens a promising avenue for future research, focusing on developing guidance mechanisms that ensure physical plausibility without compromising motion authenticity.
Acknowledgement
This work was supported by National Natural Science Foundation of China (No.62406195, No.62303319), Shanghai Local College Capacity Building Program (23010503100), ShanghaiTech AI4S Initiative SHTAI4S202404, HPC Platform of ShanghaiTech University, MoE Key Laboratory of Intelligent Perception and Human-Machine Collaboration (ShanghaiTech University) and Shanghai Engineering Research Center of Intelligent Vision and Imaging.
Supplementary Material
6 Algorithm derivation
$\frac{d\mathbf{x}_{t}}{dt}=\mathbf{x}_{1}-(1-\sigma_{\text{min}})\mathbf{x}_{0},$ (19)
In the marginal velocity formula, we obtain:
$=\mathbb{E}[\mathbf{x}_{1}|\mathbf{x}_{t},c]-(1-\sigma_{\text{min}})\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t},c].$ (20)
To simplify notations, here we omit the small amount of noise $\sigma_{\text{min}}$.
$\mathbf{v}_{\theta}(\mathbf{x}_{t},t,c)=\frac{\hat{\mathbf{x}}_{1}-\mathbf{x}_{t}}{1-t}.~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}$ (22)
Substitute Eq. 22 into the following equation:
$\mathbf{x}_{t^{\prime}}$ $=\mathbf{x}_{t}-(t-t^{\prime})\mathbf{v}_{\theta}(\mathbf{x}_{t},t,c)$
$=\mathbf{x}_{t}-(t-t^{\prime})\frac{\hat{\mathbf{x}}_{1}-\mathbf{x}_{t}}{1-t}$
$=\frac{(1-t)\mathbf{x}_{t}+(t-t^{\prime})\mathbf{x}_{t}-(t-t^{\prime})\hat{\mathbf{x}}_{1}}{1-t}$
$=\frac{1-t^{\prime}}{1-t}\mathbf{x}_{t}+\frac{t^{\prime}-t}{1-t}\hat{\mathbf{x}}_{1}.$ (23)
$\hat{\mathbf{x}}_{1}\leftarrow\mathbf{x}_{t_{n}}+(1-t_{n})~{}\mathbf{v}_{\theta}(\mathbf{x}_{t_{n}},t_{n},\mathbf{c}),$
$\mathbf{x}^{\prime}_{t_{n+1}}\leftarrow\frac{1-t_{n+1}}{1-t_{n}}\mathbf{x}_{t_{n}}+\frac{t_{n+1}-t_{n}}{1-t_{n}}~{}\hat{\mathbf{x}}_{1}.$
Algorithm 2 Euler Sampling algorithm with physical constraint guidance. ($\mathbf{v}$-prediction)
3:  Sample $\mathbf{x}_{0}$ from the action distribution
4:  for $n=1,2,...,N-1$ do
5:     # Estimate $\hat{\mathbf{x}}_{1}$ (Eq. 13)
6:     $\hat{\mathbf{x}}_{1}\leftarrow\mathbf{x}_{t_{n}}+(1-t_{n})~{}\mathbf{v}_{\theta}(\mathbf{x}_{t_{n}},t_{n},\mathbf{c})$
7:     # Flow maching $\mathbf{v}$-prediction sampling (Eq. 11)
9:     # Physical constraint guidance (Eq. 12)
11:  end for
7 Parameter analysis of guidance strength
We conduct a parameter analysis of guidance strength in Tab. 7.1. The experiment results show that as the guiding strength increases, the degree of penetration between actors and reactors decreases significantly, while FID increases slightly. This is because the ground truth itself has a certain degree of penetration. Thus, this task requires our new metrics and FID to collaborate in evaluating the quality of the generated results. When the guidance strength increases to a certain extent, the decrease in penetration degree is no longer significant. Therefore, we ultimately choose $\lambda_{\text{pene}}=10$. Our method achieves the lowest penetration level while maintaining the best generation quality.

### v3

@@ 4 Experiments
Our experiment setting of human action-reaction synthesis is online and unconstrained as in  [72] for its significant potential for practical applications. Online represents real-time reaction generation where future motions of the actor are not visible to the reactor, and the opposite is offline to relax the synchronicity. Unconstrained means that the intention of the actor is invisible to the reactor. To demonstrate the universality of our flow matching and physical guidance methods, we also conducted offline setting experiments.
@@
ReGenNet [72] $\underline{11.00^{\pm 0.74}}$ $\mathbf{0.749^{\pm 0.0002}}$ $13.80^{\pm 0.16}$ $\underline{22.90^{\pm 0.14}}$
@@
ReGenNet [72] $\underline{13.76^{\pm 4.78}}$ $0.601^{\pm 0.0040}$ $6.35^{\pm 0.24}$ $\underline{12.02^{\pm 0.33}}$
@@
[…]
Guidance w. $\mathcal{L}_{\text{pene}}$ $8.073^{\pm 0.1981}$ $0.741^{\pm 0.0002}$ $13.713^{\pm 0.1079}$ $24.077^{\pm 0.1370}$
4.1 Experiment setup
Datasets. We evaluate our model on NTU120-AS, Chi3D-AS and InterHuman-AS datasets with SMPL-X [49] body models and actor-reactor annotations as in [50].They contains 8118, 373 and 6022 human interaction sequences, respectively. “AS”[72] represents that they are an extended version of the original dataset [18, 42, 64, 39], which adds annotations to distinguish actor-reactor order of each interaction sequence and SMPL-X body models for more detailed representations. We adopt the 6D rotation representation [82] in all our experiments.
4.2 Comparison to baselines
For the NTU120-AS dataset in Tab. 2 and Chi3D-AS dataset in Tab. 3, our proposed ARFlow notably outperforms baselines in terms of the FID metric, demonstrating that our method better models the mapping between the action and the reaction distribution. Our method achieves the best FID and multi-modality, second best for the action recognition accuracy and diversity on the NTU120-AS dataset and the best FID and multi-modality, second best for the diversity on the Chi3D-AS dataset. For a fair comparison, we use the pre-trained action recognition model in [72], so our action recognition accuracy is very close to the results of [72]. Given the restricted size of the Chi3D-AS test set, some fluctuations in the experimental results are to be expected. The results of InterHuman-AS dataset in Appendix B.1 show our method also yields the best results compared to baselines. As for the offline setting, we replace the Transformer decoder units equipped with attention masks with an 8-layer Transformer encoder architecture just like ReGenNet [72]. As demonstrated in Tab. B.2, our model also achieves superior performance on most of the metrics.
4.3 Physical Constraint Guidance
Qualitative results. In Fig. 4, for the same action sequences sampled from the test sets of NTU120-AS datasets, visualization results demonstrate that both MDM and ReGenNet produce varying degrees of penetration between the actor and the reactor. In contrast, our method not only produces more physically plausible reactions, i.e., mitigate penetration problems, but also more responsive reactions, i.e., move backward when kicked, due to stronger modeling ability for causal relationship between actions and reactions. For more visualizations and videos of the generated human reactions, please refer to the supplementary materials.
Quantitative results. In Tab. 1, our ARFlow with physical constraint guidance achieves the lowest Intersection Volume, Intersection Frequency and FID than other baselines, which shows that our method achieves the lowest level of penetration while ensuring the highest generation quality. In addition, the performance is further improved through our ARFlow sampling with improved physical constraint guidance, which demonstrates the effectiveness of our guidance method. Although our guidance method effectively suppresses penetration, it also leads to a slight increase in FID, as FID only measures the similarity between generated results and the ground truth distribution and the dataset itself exists a certain degree of penetration.
4.4 Ablation Study
Network Prediction. As depicted in Tab. 3.1, a straightforward and effective strategy is to estimate clean body poses directly through a neural network, i.e., $\mathbf{x}_{1}$-prediction. We compared it with $\mathbf{v}$-prediction and the results are listed on the Prediction setting in Tab. 4 and  Tab. B.3. Obviously, $\mathbf{x}_{1}$-prediction has demonstrated superior performance across both settings. The reason we analyze it is that the geometric losses to regularize the generative network during the training phase directly acts on the predicted clean body poses, while $\mathbf{v}$-prediction requires using the predicted vector field to estimate the clean poses, so the models trained by $\mathbf{x}_{1}$-prediction will achieve stronger overall performance.
Physical constraint guidance. As we discussed earlier, the results in Tab. 1 indicate that our physical constraint guidance can effectively suppress penetration occurrences, and it may also lead to a decrease in some other metrics shown in Tab. 4 and Tab. B.3. We also provide a qualitative comparison of the effects before and after using our physical constraint guidance in Tab. H.1. The qualitative and quantitative results demonstrated that our method achieves the lowest penetration level while maintaining the best quality of generated reactions.
Number of Euler sampling timesteps. We present comprehensive evaluation results in both online and offline scenarios, with varying Euler sampling intervals (2, 5, 10 and 100 timesteps), including the latency of reaction generation per frame on online settings and overall latency on offline settings. The experimental results, as detailed in Tab. 4 and Tab. B.3, suggest that the 5-timestep Euler sampling consistently achieves optimal performance, demonstrating superior FID scores while maintaining low latency across both evaluation settings. Thus, we adopt the 5-timestep inference as the standard configuration like  [72] for all the experimental results reported in this study.
5 Conclusion
In this work, we have presented Action-Reaction Flow Matching (ARFlow), a novel framework for human action-reaction synthesis that addresses the limitations of existing diffusion-based approaches. By establishing direct action-to-reaction mappings through flow matching, ARFlow eliminates the need for complex conditional mechanisms and computationally expensive noise-to-data transformations. ARFlow involves a reprojection sampling algorithm with physical constraint guidance to enable efficient, physically plausible motion generation while preventing body penetration artifacts. Extensive evaluations on the NTU120, Chi3D and InterHuman datasets demonstrate that ARFlow excels over existing methods, showing superior performance in terms of Fréchet Inception Distance and motion diversity. Additionally, it significantly reduces body collisions, as evidenced by our new Intersection Volume and Intersection Frequency metrics.
Limitations. Although we attempt to use a reprojection method to address the issue of manifold distortions—deviations from the natural motion distribution established by flow matching, this problem still exists and our penetration loss function is just a simple design. Addressing these challenge opens a promising avenue for future theoretical research, focusing on developing guidance mechanisms that ensure physical plausibility without compromising motion authenticity.
Acknowledgement
This work was supported by National Natural Science Foundation of China (No.62406195, No.62303319), Shanghai Local College Capacity Building Program (23010503100), ShanghaiTech AI4S Initiative SHTAI4S202404, HPC Platform of ShanghaiTech University, Core Facility Platform of Computer Science and Communication of ShanghaiTech University, and MoE Key Laboratory of Intelligent Perception and Human-Machine Collaboration (ShanghaiTech University) and Shanghai Engineering Research Center of Intelligent Vision and Imaging.
Supplementary Materials
[…]
Appendix
Appendix A Algorithm derivation
$\mathbf{x}_{t}=t\mathbf{x}_{1}+[1-(1-\sigma_{\text{min}})t]\mathbf{x}_{0},~{}~{}~{}~{}~{}~{}~{}~{}~{}$ (20)
$\frac{d\mathbf{x}_{t}}{dt}=\mathbf{x}_{1}-(1-\sigma_{\text{min}})\mathbf{x}_{0},~{}~{}~{}~{}~{}~{}~{}$ (21)
In the marginal velocity formula, we obtain:
$=\mathbb{E}[\mathbf{x}_{1}|\mathbf{x}_{t},c]-(1-\sigma_{\text{min}})\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t},c].~{}~{}~{}~{}$ (22)
Substitute Eq. 23 into the following equation:
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\mathbf{x}_{t^{\prime}}$ $=\mathbf{x}_{t}-(t-t^{\prime})\mathbf{v}_{\theta}(\mathbf{x}_{t},t,c)$
$=\mathbf{x}_{t}-(t-t^{\prime})\frac{\hat{\mathbf{x}}_{1}-(1-\sigma_{\text{min}})\mathbf{x}_{t}}{1-(1-\sigma_{\text{min}})t}$
$=\frac{1-(1-\sigma_{\text{min}})t^{\prime}}{1-(1-\sigma_{\text{min}})t}\mathbf{x}_{t}+\frac{t^{\prime}-t}{1-(1-\sigma_{\text{min}})t}\hat{\mathbf{x}}_{1}.$ (24)
$\hat{\mathbf{x}}_{1}\leftarrow(1-\sigma_{\text{min}})\mathbf{x}_{t_{n}}+(1-(1-\sigma_{\text{min}})t_{n})~{}\mathbf{v}_{\theta}(\mathbf{x}_{t_{n}},t_{n},\mathbf{c}),$ (25)
$\mathbf{x}^{\prime}_{t_{n+1}}\leftarrow\frac{1-(1-\sigma_{\text{min}})t_{n+1}}{1-(1-\sigma_{\text{min}})t_{n}}\mathbf{x}_{t_{n}}+\frac{t_{n+1}-t_{n}}{1-(1-\sigma_{\text{min}})t_{n}}~{}\hat{\mathbf{x}}_{1}.$ (26)
Algorithm 3 Sampling algorithm with vanilla guidance of physical constraints. ($\mathbf{v}$-prediction)
3:  Sample $\mathbf{x}_{0}$ from the action distribution
4:  for $n=1,2,...,N-1$ do
5:     # Estimate $\hat{\mathbf{x}}_{1}$ (Eq. 25)
6:     $\hat{\mathbf{x}}_{1}\leftarrow(1-\sigma_{\text{min}})\mathbf{x}_{t_{n}}+(1-(1-\sigma_{\text{min}})t_{n})~{}\mathbf{v}_{\theta}(\mathbf{x}_{t_{n}},t_{n},\mathbf{c})$
7:     # Flow maching $\mathbf{v}$-prediction sampling (Eq. 8)
9:     # Physical constraint guidance
11:  end for
Appendix B Experimental results on the InterHuman-AS dataset and offline settings
For the text-conditioned setting, we adopt T2M [20], MDM [62], MDM-GRU [62], RAIG [60] and InterGen [39] as baselines. Tab. B.1 shows our method also yields the best results compared to baselines. To demonstrate the universality of our ARFlow and physical guidance methods, we also conducted offline setting experiments in Tab.B.2 and Tab.B.3.
[…]
Real $0.722^{\pm 0.004}$ $0.002^{\pm 0.0002}$ $3.503^{\pm 0.011}$ $5.390^{\pm 0.058}$
T2M [20] $0.224^{\pm 0.003}$ $32.482^{\pm 0.0975}$ $7.299^{\pm 0.016}$ $4.350^{\pm 0.073}$ $0.719^{\pm 0.041}$
MDM [62] $0.370^{\pm 0.006}$ $3.397^{\pm 0.0352}$ $8.640^{\pm 0.065}$ $4.780^{\pm 0.117}$ $2.288^{\pm 0.039}$
MDM-GRU [62] $0.328^{\pm 0.012}$ $6.397^{\pm 0.2140}$ $8.884^{\pm 0.040}$ $4.851^{\pm 0.081}$ $2.076^{\pm 0.040}$
RAIG [60] $0.363^{\pm 0.008}$ $2.915^{\pm 0.0292}$ $7.294^{\pm 0.027}$ $4.736^{\pm 0.099}$ $2.203^{\pm 0.049}$
InterGen [39] $0.374^{\pm 0.005}$ $13.237^{\pm 0.0352}$ $10.929^{\pm 0.026}$ $4.376^{\pm 0.042}$ $\mathbf{2.793^{\pm 0.014}}$
ReGenNet [72] $0.407^{\pm 0.003}$ $2.265^{\pm 0.0969}$ $6.860^{\pm 0.0040}$ $5.214^{\pm 0.139}$ $2.391^{\pm 0.023}$
ARFlow $\mathbf{0.434^{\pm 0.003}}$ $\mathbf{1.637^{\pm 0.0413}}$ $\mathbf{3.949^{\pm 0.0042}}$ $\mathbf{5.259^{\pm 0.117}}$ $2.502^{\pm 0.021}$
Table B.1: Comparison to state-of-the-arts on the online, unconstrained setting for human action-reaction synthesis on the InterHuman-AS dataset. $\rightarrow$ denotes that the result closer to Real is better, and $\pm$ represents 95% confidence interval. We highlight the best result in Bold.
@@ Appendix B Experimental results on the InterHuman-AS dataset and offline settings
ReGenNet [72] $\underline{6.19^{\pm 0.33}}$ $0.772^{\pm 0.0003}$ $14.03^{\pm 0.09}$ $\underline{25.21^{\pm 0.34}}$
@@
Table B.2: Results on the offline, unconstrained setting on NTU120-AS. We highlight the best result in Bold and the second best in underline.
@@
Guidance w. $\mathcal{L}_{\text{pene}}$ $5.048^{\pm 0.1167}$ $0.750^{\pm 0.0002}$ $13.838^{\pm 0.0893}$ $25.048^{\pm 0.1595}$
@@
Table B.3: Ablation studies on the offline, unconstrained setting on the NTU120-AS dataset. Bold indicates the best result in our method.
Appendix C Influence of sampling randomness
As depicted in Tab. C.1, although stochastic sampling increases the diversity of generated reaction motions, it sometimes has some impact on the quality of the sample due to its stochastic nature.
[…]
Randomness $\beta$ 0.05 $13.821^{\pm 0.2895}$ $0.709^{\pm 0.0003}$ $14.002^{\pm 0.1055}$ $24.269^{\pm 0.1363}$
0.02 $8.060^{\pm 0.1517}$ $0.728^{\pm 0.0002}$ $13.928^{\pm 0.1076}$ $24.161^{\pm 0.1512}$
0.01 $7.671^{\pm 0.1357}$ $0.728^{\pm 0.0002}$ $13.895^{\pm 0.1080}$ $24.114^{\pm 0.1486}$
ARFlow 0 $7.894^{\pm 0.1814}$ $\mathbf{0.743^{\pm 0.0002}}$ $\mathbf{13.599^{\pm 0.1005}}$ $24.105^{\pm 0.1310}$